1.2.13 Question Help Find the general solution of the system whose augmented matrix is given below 1 -2 0 1 04 0 1 0 0-3 1 0 0 0 1 9 4 Select the correct choice below and, if necessary, fill in the answer boxes to complete your answer The system is inconsistent. x2 is free X3 is free x3 is free X4 is free x5 is free x5 is free
Answer #1

C.

x_1-x_2-x_4=-4\\ x_2-3x_5=1\\ x_4+9x_5=4\\ x_3=free,\ x_5=free\ say\ c\\ 2nd\ equation\ gives\ x_2=1+3c\\ 3rd\ equation\ gives\ x_4=4-9c\\ 1st\ equation\ gives\ x_1=-4+x_2+x_4\\ i.e.\ x_1=-4+(1+3c)+(4-9c)=-4+1+3c+4-9c\\ x_1=1-6c

x_1=1-6x_5,\ x_2=1+3x_5,\ x_3=free,\ x_4=4-9x_5,\ x_5=free